OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P?

OpenStudy (anonymous):

(x – 2)2 + (y + 1)2 = 64 (x + 2)2 + (y – 1)2 = 64 (x – 2)2 + (y + 1)2 = 8 (x + 2)2 + (y – 1)2 = 8

OpenStudy (anonymous):

option a

OpenStudy (anonymous):

I got 7 more question

OpenStudy (anonymous):

okay but i hope you know how to do it...

OpenStudy (anonymous):

post them and il see if i can help